# Full-waveform Approximation of Finite-Sized Acoustic Apertures: Forward and Adjoint Wavefields

### v20

@@ 6.2. Time-Reversal Operator
Here, $\mathcal{A}_{r}^{T}$ denotes the time-reversal operator associated with receiver $r$, which acts on $y_{r,e}^{\{N,D\}}$ and produces $p_{e,r}^{T}$, the time-reversed interior wavefield corresponding to emitter $e$ and receiver $r$. The time-reversal operator, as defined in Eq. (81), together with the first and second lines of Eq. (82), represents the time-reversed, interior-field counterparts of the monopole and dipole integral formulations in Eqs. (22) and (24), respectively. This operator coincides with the adjoint operator defined by Eqs. (75) and (76), along with the Cauchy initial conditions in Eq. (77) and the time-reversed regularized source defined in Eq. (78), differing only by a constant factor.
@@ 7.1. Discretized Algorithm
A bar notation is used to denote quantities in the fully discretized domain. The discretization of the wave equation system (30) on a grid staggered in both space and time is outlined in Algorithm 1.
@@
In Algorithm 1, $\Lambda^{\zeta}=e^{-\alpha^{\zeta}\Delta t/2}$ denotes the direction-dependent Perfectly Matched Layer (PML) operator, where $\alpha^{\zeta}$ is the virtual absorption coefficient along the Cartesian coordinate $\zeta$ [1, 2, 3]. Furthermore, $\bar{\mathcal{S}}_{m}^{\zeta}=(1/d)\,\bar{\mathcal{S}}_{m}$ is fixed for all directions $\zeta$, where $\bar{\mathcal{S}}_{m}$ is obtained via a discretization of either Eq. (34) or Eq. (36), depending on whether a monopole or a monopole-like dipole source is being modeled, respectively. Note that, in contrast to the velocity vector $\bar{\bm{u}}$, the acoustic density $\bar{\bm{\rho}}$ is a scalar field; however, it is virtually treated as a vector-valued quantity to ensure consistency with the direction-dependent PML formulation [1, 2, 3]. As discussed above, the assumptions underlying Eq. (36) are not strictly valid; nonetheless, it will be employed as a benchmark to illustrate the effectiveness of the force source formulation introduced in Eq. (37).
Similarly, $\bar{\mathcal{S}}_{f}^{\zeta}$ denotes the discretized form of the $\zeta$-component of the vector-valued force source defined in Eq. (37).
@@ 7.4. Discretized Mass and Force Sources
Recall from Section 4.3.2 that using Eq. (84) to model a monopole-like dipole surface source in terms of pressure imposes far-field and omnidirectionality assumptions, which may not be valid in general.
@@ 8. Numerical Results
The accuracy of the full-waveform approach was evaluated by comparing its approximated wavefields with analytical solutions, which served as benchmarks. Specifically, the action of the causal Green’s function on a source confined to a single point was computed analytically using a frequency-domain variant of the primary Green’s formula (12), where the 3D Green’s function acts on a point source. For a source located at a single point, the spatial integral in Eq. (12) is omitted.
@@ 8.1. Regularized Source $\mathcal{S}$ Confined to a Single Point
When the volumetric source $s$, or its regularized version $\mathcal{S}$, is confined to a single point, the spatial integral in the primary Green’s formula (12) and its equivalent frequency-domain expression is eliminated. The resulting formula then describes the action of the causal Green’s function on a source localized at a single spatial point. This section evaluates the accuracy of the full-waveform approximation in reproducing this analytical action, focusing on comparisons with the exact solution obtained from the primary Green’s formula (12). [^4] As discussed at the beginning of Section 2, the solution of the wave equation (1) assumes the source is square-integrable in space and time. However, this condition is not satisfied for a source defined at a single point. As a result, for a point source, all quantities included in the wave-equation system 30 are not physical quantities. Specifically, for this simplified and nonphysical numerical experiment using a $d$-dimensional radiation source $s$ at a single point, these quantities will have units of $\text{m}^{-\text{d}}$ multiplied by the units of physical quantities associated with a finite-sized source. However, integrating over a set of sampled points distributed across a finite volume of the source yields a wavefield with physically meaningful units.
@@
For a regularized source localized at an arbitrary single point $\bm{x}_{0}$, the regularized mass source satisfies Eq. (31), and is discretized as
@@ 8.1.2. Results
Furthermore, Figures 2 and 2 depict the amplitude and phase, respectively, of the wavefield at a single frequency of 1 MHz, evaluated across all receiver positions. These results further confirm the agreement between the analytical solution and the full-waveform approximation in approximating the wavefield generated by the radiation source $s$ and its regularized counterpart $\mathcal{S}$, respectively, both defined at a single point.
@@ 8.2.2. Results
Figure 6 presents the relative error (RE) between the full-waveform and analytical wavefields for all 64 receiver positions. Receiver indices are ordered sequentially by varying $\theta$, $\varphi$, and $r$. Within each group of four consecutive receivers (e.g., 1–4, 5–8, …, 61–64), the parameters $\varphi$ and $r$ are held constant, and only $\theta$ varies. Due to the emitter disk’s rotational symmetry, these intra-group positions yield nearly identical wavefield responses, resulting in closely matched RE values. Notably, receiver sets such as 1–4, 17–20, 33–36, and 49–52—corresponding to $\varphi=0$—are spatially redundant and produce identical RE values. These redundant indices are retained in the plot for completeness and consistency with earlier figures.
@@ 8.3.1. Experiment
The emitter disk, shown in yellow in Figure 3, is excited by a source pulse $p$, illustrated in the time domain in Figure 3. Under the soft-baffle assumption, the dipole integral formula (24) is derived, representing the integral of the normal derivatives of the causal Green’s function acting on a source $p$ confined to the surface. For the analytic approximation, this formula is reformulated as an integral involving the obliquity-corrected causal Green’s functions acting on the surface source, decomposed into far-field and near-field components, as shown in the second line of Eq. (24). The analytic evaluation of this integral formula is performed using the Field II toolbox.
@@ 9. Discussion and Conclusion
Approaches for solving the acoustic wave equation can be broadly categorized into analytic and full-waveform methods. Analytic methods are suitable for homogeneous or weakly heterogeneous (smoothly varying) media, as they can account for refraction effects and singly scattered waves [62, 63, 66, 67]. In contrast, full-waveform approaches are well-suited for handling complex heterogeneities, sharp transitions, and higher-order scattering phenomena in acoustic media [3].
In some Geophysical applications, acoustic waves propagate with wavelengths significantly larger than the size of the acoustic aperture. For such cases, transducers are often approximated as single points [11]. However, this assumption is not applicable to many biomedical applications where high frequencies are utilized [59, 60, 68]. Therefore, to achieve accurate approximations of acoustic waves using full-waveform approaches, particularly in terms of amplitude, it is essential to account for the finite-size effects of acoustic apertures.
This study explored the equivalence between analytic and full-waveform approaches. Specifically, we demonstrated how scalar-valued mass sources and vector-valued force sources should be defined, discretized, and incorporated into the full-waveform approach (Algorithm 1) to ensure solutions align with their associated analytic formulas.
To achieve this, an equivalence was first established between the analytical action of the causal Green’s function on a volumetric radiation source $s$ confined to a single point, and a full-waveform approximation employing a regularized variant of this source.
@@
To better illustrate the practical impact of our study, we provide an example. An analytic approximation of a time-reversed variant of the dipole integral formula (the second line in Eq. (24)) has been widely used as the back-projection step in photoacoustic tomography for acoustically homogeneous media, under the term universal backprojection [59, 60, 68]. Our derived adjoint formula, under the practical assumption of Dirichlet-type measured data on the boundary, coincides with this back-projection operator in the homogeneous case. The practical implications of this study may grow significantly in the emerging context of therapeutic ultrasound optimization, where errors in wavefield reconstruction can raise safety concerns.
Acknowledgment
This study was conducted at the Department of Biomedical Engineering, School of Electrical and Computer Engineering, University College of Engineering, University of Tehran. The work presented in Section 8.1 was supported by the UK EPSRC Grant under Project Reference: EP/T014369/1, and carried out independently of the author’s roles at the Department of Medical Physics & Biomedical Engineering, University College London.

### v21

@@ 6.2. Time-Reversal Operator
Here, $\mathcal{A}_{r}^{T}$ denotes the time-reversal operator associated with receiver $r$, which acts on $y_{r,e}^{\{N,D\}}$ and produces $p_{e,r}^{T}$, the backprojected interior wavefield corresponding to emitter $e$ and receiver $r$. The time-reversal operator, as defined in Eq. (88), together with the first and second lines of Eq. (89), represents the time-reversed, interior-field counterparts of the monopole and dipole integral formulations in Eqs. (22) and (24), respectively. Taking the point discussed in Remark 2 into account, a regularized version of this operator coincides with the adjoint operator defined by Eqs. (78) and (79), together with the Cauchy initial conditions in Eq. (80) and the time-reversed regularized source in Eq. (81), differing only by a constant factor.
@@ 7.1. Discretized Algorithm
A bar notation is used to denote fields in the fully discretized domain. The discretization of the wave equation system (30) on a grid staggered in both space and time is outlined in Algorithm 1.
@@
In Algorithm 1, $\Lambda^{\zeta}=e^{-\alpha^{\zeta}\Delta t/2}$ denotes the direction-dependent Perfectly Matched Layer (PML) operator, where $\alpha^{\zeta}$ is the virtual absorption coefficient of the PML layer along the Cartesian coordinate $\zeta$ [1, 2, 3]. To accommodate this directional dependency, the scalar mass source is decomposed along the Cartesian axes as $\bar{\mathcal{S}}_{m}^{\zeta}=(1/d)\,\bar{\mathcal{S}}_{m}$, which remains identical in all directions $\zeta$ under the assumption of isotropy. Here, $\bar{\mathcal{S}}_{m}$ is obtained by discretizing either Eq. (34) or Eq. (36), depending on whether a monopole or a monopole-like dipole source is being modeled, respectively.
Note that, in contrast to the velocity vector $\bar{\bm{u}}$, the acoustic density $\bar{\bm{\rho}}$ is a scalar field; however, it is virtually treated as a vector-valued field to ensure compatibility with the direction-dependent PML formulation [1, 2, 3]. As discussed earlier, the assumptions underlying Eq. (36) are not strictly valid; nevertheless, owing to its widespread use [3], it will be employed as a benchmark to demonstrate the effectiveness of the force-based source formulation introduced in Eq. (37).
In addition, $\bar{\mathcal{S}}_{f}^{\zeta}$ denotes the discretized form of the $\zeta$-component of the vector-valued force source defined in Eq. (37).
@@ 7.4. Discretized Mass and Force Sources
Recall from Section 4.3.2 that using Eq. (91) to model a monopole-like dipole surface source in terms of pressure imposes two assumptions—far-field propagation and omnidirectionality—which may not be valid in general.
@@ 8. Numerical Results
The accuracy of the full-waveform approach was evaluated by comparing its approximated wavefields with analytical solutions, which served as benchmarks. Specifically, the action of the causal Green’s function on a source confined to a single point was computed analytically using a frequency-domain variant of the primary Green’s formula (12), where the 3D Green’s function acts on a point source. For a source confined to a single point, the spatial integral in Eq. (12) is omitted.
@@ 8.1. Regularized Source $\mathcal{S}$ Confined to a Single Point
When the volumetric source $s$, or its regularized version $\mathcal{S}$, is confined to a single point, the spatial integral in the primary Green’s formula (12) and its equivalent frequency-domain expression is eliminated. The resulting formula thus describes the action of the causal Green’s function on a nonphysical source confined to a single spatial point. This section evaluates the accuracy of the full-waveform approximation in reproducing this analytical action, focusing on comparisons with the exact solution obtained from the primary Green’s formula (12). [^5] For this simplified and nonphysical numerical experiment using a $d$-dimensional radiation source $s$ confined to a single point, the quantities will have units of $\text{m}^{-\text{d}}$ multiplied by the units of physical quantities associated with a finite-sized source. Integrating over a set of sampled points distributed across a finite volume of the source then yields a wavefield with physically meaningful units.
@@
For a regularized source confined to an arbitrary single point $\bm{x}_{0}$, the regularized mass source satisfies Eq. (31), and is discretized as
@@ 8.1.2. Results
Furthermore, Figures 2 and 2 depict the amplitude and phase, respectively, of the wavefield at a single frequency of 1 MHz, evaluated across all receiver positions. These results further confirm the agreement between the analytical solution using the radiation source $s$, confined to a single point, and the full-waveform approximation of its regularized counterpart $\mathcal{S}$.
@@ 8.2.2. Results
Figure 6 presents the relative error (RE) between the full-waveform and analytical wavefields for all 64 receiver positions. Receiver indices are ordered sequentially by varying $\theta$, $\varphi$, and $r$. Recall that within each group of four consecutive receivers (e.g., 1–4, 5–8, …, 61–64), the parameters $\varphi$ and $r$ are held constant, and only $\theta$ varies. Due to the emitter disk’s rotational symmetry, these intra-group positions yield nearly identical wavefield responses, resulting in closely matched RE values. Notably, receiver sets such as 1–4, 17–20, 33–36, and 49–52—corresponding to $\varphi=0$—are spatially redundant and produce identical RE values. These redundant indices are retained in the plot for completeness and consistency with earlier figures.
@@ 8.3.1. Experiment
The emitter disk, shown in yellow in Figure 3, is excited by a source pulse $p$, illustrated in the time domain in Figure 3. Under the soft-baffle assumption, the dipole integral formula (24) is derived, representing the integral of the normal derivatives of the causal Green’s function acting on a source $p$ confined to the surface. For the analytic approximation using the Field II toolbox, we reformulate the original dipole integral formula as an integral involving the obliquity-corrected causal Green’s functions acting on the surface source, decomposed into far-field and near-field components according to the second line of Eq. (24). The resulting analytic integral was evaluated numerically using the Field II toolbox.
@@ 9. Discussion and Conclusion
Approaches for solving the acoustic wave equation can be broadly categorized into analytic and full-waveform methods. Analytic methods are suitable for homogeneous or weakly heterogeneous (smoothly varying) media, as they can account for refraction effects and singly scattered waves [62, 63, 66, 67, 68]. In contrast, full-waveform approaches are well-suited for handling complex heterogeneities, sharp transitions, and higher-order scattering phenomena in acoustic media [1, 2, 3].
In certain geophysical applications, acoustic wavelengths are significantly larger than the size of the acoustic aperture, allowing transducers to be approximated as point-like sources or receivers [11]. This assumption, however, does not hold in many biomedical imaging scenarios, where much higher operating frequencies are employed [59, 60, 69]. Consequently, accurate full-waveform modeling—particularly when amplitude fidelity is required—necessitates accounting for the finite-size effects of acoustic apertures.
This study investigated the equivalence between analytic and full-waveform formulations. In particular, we established how scalar-valued mass sources and vector-valued force sources must be defined, discretized, and incorporated into the full-waveform framework (Algorithm 1) to ensure that the resulting solutions are consistent with their corresponding analytical expressions.
To this end, we first demonstrated an equivalence between the analytical action of the causal Green’s function on a volumetric radiation source $s$ confined to a single point and its full-waveform counterpart obtained using a regularized approximation of this source.